# Taming OOD Actions for Offline Reinforcement Learning: An Advantage-Based Approach

### v3

@@ C.7 Auxiliary Model Pretraining
Architecture 4-layer MLP 256 neurons per layer 5-layer MLP 256 neurons per layer
@@
C.8 Dataset Construction for Offline RL in PointMaze
@@
To construct the offline dataset, we collect trajectories across multiple runs of online SAC [12] training. We employed a staged sampling strategy during online training (Algorithm 3). This strategy involved periodically performing trajectory rollouts using the current policy, allowing us to collect a diverse set of behaviors as the policy evolved throughout training. This collection process spanned 10 independent runs of online SAC training, each for up to 250 000 steps.
@@
In total, the collected offline dataset comprises 853 trajectories. As detailed in Table 5, the dataset exhibits a distribution heavily skewed towards suboptimal behavior, consistent with our collection strategy. Specifically, it contains no optimal trajectories, a very small number of near-optimal paths (3.4%), and a modest proportion (15.7%) of competitive paths, which we consider to be acceptably successful for the task. The vast majority of the dataset (80.9%) consists of trajectories categorized as Sub-Optimal based on their length, defined as exceeding 300 steps.

### v4

@@ C.6 Auxiliary Model Pretraining
Architecture
4-layer MLP
256 neurons per layer
5-layer MLP
256 neurons per layer
@@
C.7 Dataset Construction for Offline RL in PointMaze
@@
To construct the offline dataset, we collect trajectories across multiple runs of online SAC (Haarnoja et al., 2018) training. We employed a staged sampling strategy during online training (Algorithm 2). This strategy involved periodically performing trajectory rollouts using the current policy, allowing us to collect a diverse set of behaviors as the policy evolved throughout training. This collection process spanned 10 independent runs of online SAC training, each for up to 250 000 steps.
We manually collected 853 successful yet sub-optimal trajectories to construct the offline training dataset, each shorter than 1000 steps but significantly longer than the shortest length of approximately 142 steps, which corresponds to a straight-line path from the start to the goal. As shown in the last subfigure of Figure˜4, the offline dataset includes no optimal trajectories (length << 175), only 3.4% near-optimal (length << 225), and 19.1% competitive trajectories (length << 300). These trajectories fall into three distinct modes—Left, Middle, and Right routes—depending on which corridor the agent takes to bypass the obstacles (see corresponding subfigures in Figure˜4). The dataset the bar plot of Figure 4 comprises 391 391 Q-learning-style tuples of the form $(\bm{s},\bm{a},\bm{s^{\prime}},r,\text{done})$.
@@
As detailed in Table 5, the dataset exhibits a distribution heavily skewed towards suboptimal behavior, consistent with our collection strategy. Specifically, it contains no optimal trajectories, a very small number of near-optimal paths (3.4%), and a modest proportion (15.7%) of competitive paths, which we consider to be acceptably successful for the task. The vast majority of the dataset (80.9%) consists of trajectories categorized as Sub-Optimal based on their length, defined as exceeding 300 steps.
@@
Appendix D Declaration
I declare that Large Language Models (LLMs) were used solely for language polishing in this paper. No other usage of LLMs was involved.